Inflation took its toll, and the merchant had to mark the suit up 30 percent so it would sell for $156. What was the original price of the suit?
Answer by Theo(13342)

x represents the original price.
1.3 * x = 156
x = 156 / 1.3 = 120
original price was 120.
mrked up prive was 120 + .3 * 120 = 120 + 36 = 156.
solution is that the original price was 120.